If C=100+0.75Y

T=50+0.5Y

X=200

M=50+0.25Y

G=150

Determine the equilibrium level of output


1
Expert's answer
2021-05-03T10:51:09-0400

"Y=C(Y-T)+X+G-M;"

"Y=100+0.75\u00d7(Y-(50+0.5Y))+200+150-(50+0.25Y);"

"Y=100+0.75\u00d7(0.5Y-50)+200+150-50-0.25Y;"

"Y=100+0.375Y-37.5+300-0.25Y;"

"Y=362.5+0.125Y;"

"0.875Y=362.5;"

"Y=414.3;"